What is the smallest possible value of the principal quantum number n for an s electron? -n=1 x 5 ?
Answer #1

Answer : n = 1 is the smallest possible value of the principal quantum number n for a s electron.

explanation:

Explanation:

We know if an electron is in the s orbital which the means that orbital angular momentum quantum number l will be 0.

And, the principal quantum number n and l is related by:

l=0,1,2,3,...(n−1)

Therefore nmin=l+1

=0+1

=1.

please upvote